For any integer "n", find i^(4n+a) when "a" takes on the following values.
3 & 2
Please show all work-- I don't understand this type of problem

Answer by richard1234(7193)

Hint: i^1 = i, i^2 = -1, i^3 = -i, i^4 = 1 and it repeats.

Also, i^(4n+a) = i^(4n)i^a = i^a

Now just plug in a = 2, 3.